Shift Summary for Evening, 6/21/99

Elliott Cheu (
Mon, 21 Jun 1999 22:45:33 -0700

Shift Summary for EVENING shift, Jun. 21, 1999

On Shift: ECC, UE
Also: Hogan, Sydney and Ashkan.

Hours of beam delivered at all intensities: 7.0h
Hours of beam delivered at nominal intensity: 7.0h
Hours of E832 data (special runs not included): 6.7h
Hours of special runs (laser scans, muon runs...): 0.3h

Accounting of lost time: kfrends, changing tapes

Runs ended this shift:

Run 13680 - Laser - Last few settings with beam.
13681 - 14.1M - BEAM_832
13682 - BEAM_MU_4E12
13683 - BEAM_MU_4E12

Tapes brought to FCC this shift: QKC028-31, QKS028-31, QKK028-31

Accesses during shift: Marj and CW to look at HA#28. Problem seems to
be with shaper splitter

Overview of shift:

Mainly smooth E832 datataking at 5E12.

Description of Shift:

16:20 Beam returns while doing laser scan. Last few laser settings with

16:40 Onsplit requested QKLS01 to be mounted.

17:25 NM2SR noticed to be off. Called MCR and it was turned back on.

18:35 ktev1a filling up. Deleted files. Theo changed warning limit to
95% from 99%.

19:20 Two CsI channels: 1642 and 2310 have some problems. Ed and PSS notified.

20:30 ktevana errors. These come from crate 53. It happens about three
times during this run.

22:20 Hogan notes that L1 timing seems to have shifted. RSK suggests
adding 1ns to TeV RF signal. This is done for subsequent muon

23:45 Run 13681 ended because tapes filled up. Removed tapes and
switched to muon mode. Switched magnet polarity to POS.

23:55 ktpipe04 failed to init. Rebooted.

00:00 Start muon run 13682. Note TeV RF delayed by 1ns.

00:05 Muon run 13683 with 1ns delay removed.

00:06 Muon run ended. Added back in 1ns of TeV RF.

Items to Follow up on Next Shift

ktev2 crashed as new BEAM_MU4E12 run was ending.